In a particular double-slit experiment, the distance between the slits is 50 times the wavelength of...

  1. In a particular double-slit experiment, the distance between the slits is 50 times the wavelength of the light. Find the angle (in degrees) at which the first-order maximum occurs.
  1. 2.78
  2. 1.15
  3. 5.60
  4. 0.01
Answer #1

For nth maxima,

d sin\ \theta=n\lambda

Given ,

d = 50\lambda

And for first order maxima, n = 1

Therefore,

(50\lambda)\sin\theta=\lambda

\sin\theta=\frac {\lambda}{ 50\lambda}

\ \theta=sin^{-1}\frac {1}{ 50}

\boxed {\ \theta=1.15\degree}

Therefore, the correct option is (b).